# Optical excitation of bulk plasmons in n-doped InAsSb thin films : investigating the second viscosity in electron gas

### v1

@@ Abstract
We demonstrate that including the second viscosity of an electron gas in the hydrodynamic model allows for highly accurate modeling of the optical response of heavily doped semiconductors. In our setup, which improves resonance visibility compared to previous approaches, plasmon resonances become more distinct, allowing for detailed analysis of the underlying physics. With advanced fitting techniques based on a physics-informed cost function and a tailored optimization algorithm, we obtain close agreement between simulations and experimental data across different sample thicknesses. This enhanced resonance visibility, combined with our integrated approach, shows that key parameters such as doping level and effective electron mass can be retrieved from a single optical measurement. The spatial dispersion taken into account in the hydrodynamic framework is essential for accurately describing the optical response of plasmonic materials in this frequency range and is likely to become a standard modeling approach.
@@
The optical response of highly n-doped semi-conductors is essentially dictated by the response of the free electron gas, which is usually and effectively described using a simple Drude model[1]. This model assumes no interactions between electrons within the gas. However, since the early work of Tonks and Langmuir in the 20th century[2], it has been established that electron gases can also support plasmons – longitudinal electromagnetic waves similar to sound. Accurately modeling these waves requires moving beyond the Drude model to include electron-electron repulsion at the scale of the mean free path typically, which introduces spatial dispersion, or nonlocality, into the description.
Near the plasma frequency, even a single slab of doped semiconductors exhibits resonances. These resonances, first predicted for thin metal films by Ferrell[3] using the Drude model, can be optically excited and are now routinely employed to characterize doping levels in semiconductors[4].
Melnyk and Harrison[5] have been the very first to suggest that longitudinal waves are actually excited inside the slab and are responsible even for the Ferrel resonance. They provided a theoretical framework to describe the response of a metallic slab that is still relevant nowadays[6]. Using Maxwell’s equations and a simple hydrodynamic model for the electron gas, they predict that plasmon resonances should occur in a metallic film just above the screened plasma frequency, for which the permittivity vanishes. They even suggest that this could happen for doped semi-conductors as well. Such resonances can be considered as cavity resonances for plasmons, with some peculiar characteristics however.
Only recently has the scientific community started to realize the importance of nonlocality in doped semiconductors[7]. Recent experiments have shown resonances linked to the excitation of plasmons in semi-conductor films of 20 nm or less in thickness[8], amid a debate on the origin of the supplementary losses experienced by plasmons[9], beyond those due to friction with the lattice. Such resonances, observed in quantum wells, have also been and described using a fully quantum approach to the electron gas[10]. We underline that spatial dispersion can have an impact also below the (screened) plasma frequency for guided modes with a large wavevector, as is common in plasmonics. Therefore, investigating the effects of nonlocality is crucial for a deeper understanding and a more accurate modeling of these phenomena.
In this work, we demonstrate that cavity resonances for plasmons can be observed in doped semiconductors that are nearly an order of magnitude thicker than those previously studied[8]. In order to do so, we illuminate our sample using a high index germanium prism, in a relatively symmetrical configuration which enhances the visibility of the resonances. We provide new physical insights to better understand their characteristics. By utilizing advanced numerical fitting techniques, we show that the experimental spectra can be almost perfectly replicated using the formalism inspired by the hydrodynamic model[11]. Importantly, we show that taking into account the bulk viscosity of the electron gas results in an imaginary part of the nonlocal parameter $\beta^{2}$[12, 9, 8], and that this is necessary to reach a satisfactory agreement with experimental data.
Our results suggest that accounting for spatial dispersion can enable the accurate determination of the electron gas characteristics, such as its density, the effective mass of electrons and more, in a single optical measurement.
@@
In the first part of this paper, we review the basics of spatial dispersion in an electron gas and the characteristics of plasmons to provide a foundation for a more thorough discussion of our results. We then detail our experimental findings, demonstrating how the observed resonance positions align with predictions from the hydrodynamic model. In the third section, we explain how the use of a global optimization algorithm and an adapted cost function to fit our experimental data allows to retrieve all the model parameters. The final section is dedicated to an in-depth discussion of the overall implications of our results.
@@ I.1 Hydrodynamic model
It should be underlined that the effective polarization, in the harmonic regime (we assume a $e^{-i\omega t}$ time dependency here), is actually proportional to the displacement of the electrons (or carriers more generally) with respect to their average position. The second term is linked to the friction of the electrons on the lattice, while the last is the electric force. In the harmonic regime, the displacement of the electrons is always opposite to the electric force because there is no other restoring force. The effective polarization is thus opposite to the electric field and dominates if the frequency is low enough, making the permittivity negative.
@@
The screened plasma frequency, for which the real part of the permittivity vanishes, is given approximately by
@@
It is called the screened plasma frequency or the epsilon-near-zero (ENZ) frequency in the following because below this frequency the semiconductor can be considered as a plasmonic material, like metals in the visible range. Although many authors use ”plasma frequency” to describe the ENZ frequency, these two frequencies must be distinguished, particularly given the high susceptibility of semiconductors in the infrared range (typically, $\chi_{b}\simeq 10$).
It should also be noted that the actual plasma frequency $\omega_{p}$ cannot be measured directly. Experimental measurements always yield the screened plasma frequency $\omega_{0}$, as the response of the surrounding medium is inherently included in any optical observation.
The hydrodynamic model [13] is the simplest possible way to take into account the repulsion between electrons, by integrating a pressure term in the description of the electron gas response. This leads to the following relation between the electric field and the effective polarization:
$\frac{\partial^{2}\mathbf{P_{f}}}{\partial t^{2}}+\gamma\frac{\partial\mathbf{P_{f}}}{\partial t}-\beta^{2}\nabla(\nabla.\mathbf{P_{f}})=\varepsilon_{0}\omega_{p}^{2}\mathbf{E}.$ (4)
The electron pressure term $\beta^{2}\nabla(\nabla.\mathbf{P_{f}})$, coming from the Thomas–Fermi theory of metals, obviously includes spatial derivatives of the polarization, making the description non-local. We underline that the continuity equation can be written .$\nabla.\mathbf{P_{f}}=-\rho$ in this framework, so that this supplementary term truly appears as a force pushing electrons away from any concentration of negative charges.
@@
The parameter $\beta$ quantifies the non-local effects and is called the hydrodynamic parameter. It is usually defined as $\beta=\sqrt{3/5}\,v_{F}$ with $v_{F}=\frac{\hbar}{m^{*}}(3\pi^{2})^{1/3}n_{0}^{1/3}$, the Fermi velocity [13]. The coefficient between $\beta$ and $v_{F}$ actually depends on the frequency. However, close to the screened plasma frequency $\sqrt{3/5}$ is a very accurate assumption[12].
Whatever the medium considered, the electric field can always be decomposed as the sum of a divergence-free component (the transverse wave) and a curl-free component (the longitudinal wave)[14]. Since the divergence of the magnetic field is always null, the magnetic field always belongs to the purely transverse part of the decomposition. The two kinds of waves evolve independently from each other – in the sense that the transverse wave is never converted into a longitudinal one along its propagation in a homogeneous medium. The transverse wave is completely insensitive to spatial dispersion, it is described as a wave propagating in a local medium with a permittivity given by the Drude model. We underline generally that while classical electromagnetic theorems can be extended to spatially dispersive media, their generalization is not straightforward[15].
At an interface between the non-local medium and any local medium, the component of the current perpendicular to the interface vanishes, which is simply given in the harmonic regime by $\mathbf{P_{f}}\cdot\mathbf{n}=0$ and provides the necessary additional boundary condition[16, 17].
When light from a local medium illuminates an interface, the two types of waves (transverse and longitudinal) can be excited within the nonlocal medium. The boundary conditions determine the amplitude of these waves. More specifically, the strength of the longitudinal wave excitation depends on the direction of the electric field at the interface. At normal incidence, with the electric field purely tangential to the surface, the longitudinal wave remains unexcited, but as the incidence angle increases, the excitation of the longitudinal wave becomes stronger.
When a longitudinal wave inside a nonlocal material encounters an interface with a local material, it is reflected, but a transverse wave is also excited in the process. This phenomenon is best described using a scattering matrix formalism[11] adapted to the hydrodynamic model, which we employ here to calculate the reflectance of any given structure.
In what follows, we will use the term plasmon to refer to the longitudinal wave, though technically, a plasmon is the quantum of this wave. The transverse wave will simply be referred to as light.
The dispersion relation of the plasmon, taking into account losses due to friction with the lattice, can be written as[16]:
@@
$\omega^{2}+i\gamma\omega=\omega_{0}^{2}+\beta^{2}\mathbf{k}^{2},$ (5)
@@
An example of dispersion curve for the plasmon is shown in Figure 1. The parameters chosen here are in agreement with the experiments and data fits presented in the third section. Below the screened plasma frequency $\omega_{0}$, plasmons are evanescent, just as light. Their wavevector being essentially imaginary, it possible to define a typical penetration depth $\delta_{\mathrm{p}}$ for these waves. It is usually much shorter for plasmons than for light. All accounted for, this typical length is proportional to $n_{0}^{-1/6}m^{*-1/2}$, thus decreasing with the electronic density. In highly-doped semiconductors the free carrier density $n_{0}$ is orders of magnitude lower than in noble metals and the effective mass $m^{*}$ is also lower – this explains why $\omega_{\mathrm{p}}=\sqrt{\frac{n_{0}e^{2}}{m^{*}\varepsilon_{0}}}$ is in the infra-red range.
However, these considerations show that theoretically, the plasmon skin depth $\delta$ is at least 5 times larger or more in highly-doped semiconductors than in noble metals. This constitutes another indication that non-local effects can be expected to play a much larger role in the semiconductors response, even below the screened plasma frequency, because of its lower carrier density.
@@
Different authors have considered that the parameter $\beta$ should be taken complex and frequency-dependent, on very different physical grounds. For Halevi, the reasons lie in the microscopic model[12] that underpins the hydrodynamic model. In the Generalized Nonlocal Optical Response (GNOR) model, a supplementary diffusion term is considered[9], while it has been shown more recently that accounting for viscosity within the electron gas also results in a complex $\beta$ with a different frequency dependence[8]. To summarize, while there is consensus on the necessity of an imaginary component for $\beta$, its origin and physical meaning remain debated to this date.
I.2 Resonances in a metallic slab
In the Drude model, the very first step is to treat the currents inside the free electron gas as an effective polarization. This can be done easily because oscillating dipoles also correspond to real microscopic currents. The second step is to assume that free electrons oscillate independently under the influence of the incoming electric field.
Within this framework, when a horizontal slab containing a free electron gas is illuminated in normal incidence, the gas is free to oscillate horizontally. No resonance is associated to such an oscillation. Depending on whether the frequency is larger or smaller than the screened plasma frequency, the horizontal oscillation will allow the electric field to penetrate inside the slab or not, respectively.

### v2

@@ Abstract
We demonstrate that including the second viscosity of an electron gas in the hydrodynamic model allows for highly accurate modeling of the optical response of heavily doped semiconductors. In our setup, which improves resonance visibility compared to previous approaches, plasmon resonances become more distinct, allowing for a detailed analysis of the underlying physics. With advanced fitting techniques based on a physics-informed cost function and a tailored optimization algorithm, we obtain a close agreement between simulations and experimental data across different sample thicknesses. This enhanced resonance visibility, combined with our integrated approach, shows that key parameters such as doping level and effective electron mass, as well as the second viscosity of the electron gas, can be retrieved from a single optical measurement. The spatial dispersion taken into account in the hydrodynamic framework is essential for accurately describing the optical response of plasmonic materials in this frequency range and is likely to become a standard modeling approach.
@@
Plasma oscillations in conducting materials were first described by Tonks and Langmuir[tonks1929oscillations], who identified these longitudinal waves, similar to sound, now called plasmons, in ionized gases near the plasma frequency. Ferrell[ferrell1962plasma] later predicted that thin metal films should show resonances just above the screened plasma frequency, where the permittivity approaches zero. Melnyk and Harrison[melnyk1968resonant, melnyk1970theory] used a hydrodynamic model that included electron-electron interactions to describe the metallic response more accurately. They predicted that metal slabs should support several odd-order bulk plasmon resonances, not just the single resonance that Ferrell had predicted. These predictions were confirmed in very thin potassium films[anderegg1971optically], but for most metals the plasma frequency sits right in the middle of interband transitions, where absorption is so strong that plasmonic resonances cannot exist.
In heavily doped semiconductors, however, the plasma frequency is in the infrared range, far from interband transitions. The community first used this to develop a practical tool: the main plasmon resonance (often called the Ferrell or Epsilon-Near-Zero resonance) can be used with calibration charts to measure doping concentrations from simple optical measurements[taliercio2014brewster]. This application was developed without much concern for the theoretical predictions about higher-order resonances.
Separately, experiments on particle-on-mirror systems showed that Drude’s model[drude1900elektronentheorie] is not sufficient to accurately describe particles on mirror resonances[Ciraci2012, reynaud2018enhancing], but that the hydrodynamic model was able to do so. This has renewed interest in the hydrodynamic model, for which improvements have been proposed[Moreau2013, mortensen2014generalized]. It has also been implemented in various numerical simulation methods[Toscano2012, benedicto2015numerical, schmitt2016dgtd, kwiecien2023nonlocal, langevin2024pymoosh], making it more widely usable.
Recently, De Ceglia et al.[de2018viscoelastic] showed that very thin films of heavily n-doped semiconductors do exhibit multiple bulk plasmon resonances, confirming what the hydrodynamic model predicted decades ago. This has renewed interest in properly understanding these effects. Quantum mechanical calculations can describe these systems accurately[vasanelli2020semiconductor]; however, it seems that the hydrodynamic model offers a practical alternative that captures the main physics.
In this work, we show that bulk plasmon resonances can be observed in doped semiconductor films that are an order of magnitude thicker than those studied previously. We use a pseudo-ATR configuration where the semiconductor sits between two high-index materials, which makes the resonances much easier to observe. We find that including the bulk viscosity (or second viscosity) of the electron gas in the hydrodynamic model is necessary for accurate fits. This provides a natural explanation for why the nonlocal parameter needs an imaginary part[halevi1995hydrodynamic, mortensen2014generalized, de2018viscoelastic]. Using a global optimization algorithm combined with a physics-informed cost function, we extract material parameters —doping concentration, effective mass, and bulk viscosity— from a single optical measurement without the need for calibration charts. Our fitting procedures and analysis tools are available open-source, making this characterization technique immediately accessible to the community.
@@
In the first part of this paper, we review the basics of spatial dispersion in an electron gas and the characteristics of plasmons to provide a foundation for a thorough understanding of our results. We then detail our experimental findings, demonstrating how the observed resonance positions align with predictions from the hydrodynamic model. In the third section, we explain how the use of a global optimization algorithm and an adapted cost function to fit our experimental data allows us to retrieve all the model parameters.
@@ I.1 Drude model context
It should be underlined that the effective polarization, in the harmonic regime (we assume a $e^{-i\omega t}$ time dependency here), is actually proportional to the displacement of the electrons (or carriers more generally) with respect to their average position. The second term is linked to the friction of the electrons on the lattice, while the last is the electric force. In the harmonic regime, the displacement of the electrons is always opposite to the electric force because there is no other restoring force. The effective polarization is, thus, opposite to the electric field and dominates if the frequency is low enough, making the permittivity negative.
@@
The screened plasma frequency, for which the real part of the permittivity vanishes and which is the only frequency that can be measured directly, is given approximately by
@@
It is referred to as the screened plasma frequency or the epsilon-near-zero (ENZ) frequency because, below this frequency, the semiconductor can be considered a plasmonic material, similar to metals in the visible range. Although many authors use ”plasma frequency” to describe the ENZ frequency, these two frequencies must be distinguished, particularly given the high susceptibility of semiconductors in the infrared range (typically, $\chi_{b}\simeq 10$).
I.2 Including second viscosity
Drude’s model is not sufficient to describe and understand plasmons, as it is necessary to take into account electron-electron repulsion in the description of the electron gas. The simplest way to do so is to consider the electron gas as a Newtonian fluid and to incorporate a supplementary pressure term into the fluid equation[Ciraci2013]. After linearization, these equations are coupled with Maxwell’s equations to yield a coherent description of the optical response of the electron gas. We will follow this path here, but we will consider all viscosity terms [ash1991second, gad1995stokes, graves1999bulk, buresti2015note]. The equation governing the electron gas dynamics in this framework is
where $n$ and $m^{*}$ are the electron density and electron effective mass, respectively; $\mathbf{v}$ is the electron velocity, $p$ is the pressure inside the electron gas, $\mu$ is the (shear) viscosity, $\gamma$ is a damping factor, and $\xi$ is the second (or bulk) viscosity[graves1999bulk].
In the framework of the hydrodynamic model, viscous terms are usually not included[Moreau2013, Ciraci2013, mortensen2014generalized]. It recently became clear, however, that electron gases do exhibit a viscous behavior[bandurin2016negative, polini2020viscous]. The shear viscosity has been integrated into the hydrodynamic model in a recent work for the first time[de2018viscoelastic]. The second viscosity term is usually assumed to be negligible because this form of viscosity becomes relevant only when the fluid expands or contracts significantly. It is typically neglected when deriving the Navier-Stokes equations because most fluid flows are essentially incompressible, even if the fluid itself is very compressible, leaving shear as the only source of friction[buresti2015note]. This assumption, called the Stokes hypothesis, is so common and accurate that the bulk viscosity of a gas is challenging to estimate. Furthermore, monoatomic gases are expected to have a vanishing bulk viscosity. Electrons, also lacking internal degrees of freedom, might be expected to behave similarly. It would thus seem natural to neglect the second viscosity for electron gases.
However, when plasmons are excited, as they involve periodic compression and rarefaction of the electron density, with minimal shear deformation, bulk viscosity naturally dominates over shear viscosity. This is inspired by what occurs in classical gases, where bulk viscosity is the primary mechanism for sound attenuation and typically exceeds shear viscosity by orders of magnitude [ash1991second, cramer2012numerical]. We therefore adopt this perspective for what follows, keeping only the bulk viscosity contribution to the hydrodynamic model.
We now linearize the equations to obtain a linear response model. The velocity field $\mathbf{v}$ is considered a first order term, so that the inertial term $\mathbf{v}\cdot\nabla\mathbf{v}$ can be neglected also. The electron gas density $n$ can be decomposed into a constant term $n_{0}$ and a first-order fluctuation $n_{1}$ so that finally we can write $n_{0}e\partial_{t}\mathbf{v}\simeq\partial_{t}\,\mathbf{j}$ at the first order. This can also be done for the spatial derivatives of the bulk viscosity term. Taking these steps into account leads to
$\frac{m^{*}}{e}\,\partial_{t}\mathbf{j}=-\nabla p+\frac{m^{*}}{e}\,\gamma\,\mathbf{j}-n_{0}\,e\,\mathbf{E}-\frac{\xi}{n_{0}\,e}\,\nabla(\nabla\cdot\mathbf{j})$ (5)
The pressure term is given by quantum theories and is a power of $n$ (usually $n^{3}$ is retained [halevi1995hydrodynamic]). Whatever the exponent, the gradient of the pressure is simply proportional to the gradient of the electron gas density $n$. The nonlocal parameter $\beta$ is introduced at this stage so that
$\nabla p=-m^{*}\beta^{2}\,\nabla n.$ (6)
@@ I.2 Including second viscosity
The parameter $\beta$ quantifies the non-local effects and is called the hydrodynamic parameter. It is usually defined as $\beta=\sqrt{3/5}\,v_{F}$ with $v_{F}=\frac{\hbar}{m^{*}}(3\pi^{2})^{1/3}n_{0}^{1/3}$, the Fermi velocity [Ciraci2013]. The coefficient between $\beta$ and $v_{F}$ actually depends on the frequency. However, close to the screened plasma frequency $\sqrt{3/5}$ is a very accurate assumption [halevi1995hydrodynamic].
The continuity equation can be written
$\nabla.\mathbf{j}=-\partial_{t}(-e\,n)=e\,\partial_{t}n_{1}.$ (7)
Since $\mathbf{j}=\partial_{t}\mathbf{P_{f}}$, this equation can be integrated to yield $\nabla\cdot\mathbf{P_{f}}=en_{1}$. The pressure term can be finally written
$\nabla p=-\frac{m^{*}}{e}\beta^{2}\,\nabla(\nabla\cdot\mathbf{P_{f}}).$ (8)
This term obviously includes spatial derivatives of the polarization, making the description non-local. We underline that the continuity equation can be written $\nabla.\mathbf{P_{f}}=-\rho$ in this framework, so that this supplementary term truly appears as a force pushing electrons away from any concentration of negative charges. For the second viscosity term, we have:
$\frac{\xi}{n_{0}e}\,\nabla(\nabla\cdot\mathbf{j})=-\frac{\xi}{n_{0}e}\,\partial_{t}\nabla(\nabla\cdot\mathbf{P_{f}}),$ (9)
@@
$\left(\beta^{2}-\frac{i\omega\xi}{n_{0}m^{*}}\right)\nabla(\nabla\cdot\mathbf{P_{f}})-(\omega^{2}-i\omega\gamma)\mathbf{P_{f}}=\epsilon_{0}\omega_{p}^{2}\mathbf{E}.$ (10)
Thus, the second viscosity term introduces a frequency-dependent imaginary contribution to $\beta^{2}$. Following the notation of previous works[mortensen2014generalized] we can define a complex parameter
$\eta^{2}=\beta^{2}-i\,\omega\frac{\xi}{n_{0}m^{*}}$ (11)
which naturally incorporates bulk viscosity effects.
The inclusion of bulk viscosity thus provides a physically grounded explanation for the imaginary part of $\beta^{2}$ that has been empirically necessary to fit experimental data[Raza2013]. Unlike the GNOR[mortensen2014generalized] model’s additional diffusion currents – which raise questions about charge conservation – or shear viscosity – which produces spurious resonances[raza2011unusual] – bulk viscosity naturally yields an imaginary part proportional to $\omega$ through the physical compression of the electron gas. This suggests that the successful GNOR phenomenology in fact captures the effects of bulk viscosity, making the distinction between the GNOR and hydrodynamic models unnecessary.
I.3 Plasmons
Having established the hydrodynamic model with bulk viscosity, we now examine the nature of the waves that can propagate in such a nonlocal medium.
In a medium containing an electron gas, the electric field can be decomposed as the sum of a divergence-free component (the transverse wave) and a curl-free component (the longitudinal wave)[bhatia2012helmholtz]. Since the divergence of the magnetic field is always null, the magnetic field always belongs to the purely transverse part of the decomposition.
The two kinds of waves evolve independently from each other – in the sense that the transverse wave is never converted into a longitudinal one along its propagation in a homogeneous medium. The transverse wave is completely insensitive to spatial dispersion, it is described as a wave propagating in a local medium with a permittivity given by the Drude model so that we call it light in the following. We will continue to call the longitudinal wave plasmon, though technically, a plasmon is the quantum corresponding to this wave.
We underline generally that while classical electromagnetic theorems can be extended to spatially dispersive media, their generalization is not straightforward[sakat2021generalized].
At an interface between the non-local medium and any local medium, because the electron gas can not escape from the material in which it is contained, the component of the current perpendicular to the interface vanishes. This is simply given in the harmonic regime by $\mathbf{P_{f}}\cdot\mathbf{n}=0$ and provides the necessary additional boundary condition[Moreau2013, burda2023nonlocal] required by the hydrodynamic model to be complete.
When light from a local medium illuminates an interface, the two types of waves (transverse and longitudinal) are excited within the nonlocal medium. The boundary conditions determine the amplitude of each of these waves. More specifically, the strength of the longitudinal wave excitation depends on the direction of the electric field at the interface. At normal incidence, with the electric field purely tangential to the surface, the plasmon remains unexcited, but as the incidence angle increases, the excitation of the longitudinal wave becomes stronger.
When a longitudinal wave inside a nonlocal material encounters an interface with a local material, it is reflected, but a transverse wave is also excited in the process. This phenomenon is best described using a scattering matrix formalism[benedicto2015numerical] adapted to the hydrodynamic model, which we employ here to calculate the reflectance of any given structure, assuming a complex $\beta^{2}$ parameter.
The dispersion relation of the plasmon[Moreau2013], taking into account losses due to friction with the lattice and the second viscosity, can be written as:
@@ I.3 Plasmons
An example of dispersion curve for the plasmon is shown in Figure 1. The parameters chosen here are in agreement with the experiments and data fits presented in the following. Below the screened plasma frequency $\omega_{0}$, plasmons are evanescent, just as light. Their wavevector being essentially imaginary, it is possible to define a typical penetration depth $\delta_{\mathrm{p}}$ for these waves. It is usually much shorter for plasmons than for light. All accounted for, this typical length is proportional to $n_{0}^{-1/6}m^{*-1/2}$, thus decreasing with the electronic density. In highly-doped semiconductors the free carrier density $n_{0}$ is orders of magnitude lower than in noble metals and the effective mass $m^{*}$ is also lower – this explains why $\omega_{\mathrm{p}}=\sqrt{\frac{n_{0}e^{2}}{m^{*}\varepsilon_{0}}}$ is in the infra-red range.
However, these considerations show that theoretically, the plasmon skin depth $\delta$ is at least 5 times larger or more in highly-doped semiconductors than in noble metals. This constitutes another indication that non-local effects can be expected to play a much larger role in the semiconductors’ response, even below the screened plasma frequency.
@@
I.4 Resonances in a metallic slab